a ladder is 5m long and weighs 900N makes an angle of 70 degrees with the ground. the wall is smooth and the ladder is uniform. a person weighing 750N stands 1.5m from the ladder . calculate the coefficient of friction between the ladder and the ground

a ladder is  5m long and weighs 900N makes an angle of 70 degrees with the ground. the wall is smooth and the ladder is uniform. a person weighing 750N stands 1.5m from the ladder . calculate the coefficient of friction between the ladder and the ground

Grade:10

1 Answers

Piyush
5 years ago
Let N1 be the normal force by the wall on ladder and N2 be the normal force by ground on ladder. Let the point p of ladder which touches to ground. Balancing torque along the point p.Ie.N1*5*cos20 =900*2.5*cos70 + 750* 1.5*cos 70After calculating,N1=246.7 NBy balancing the forces, N1=force of friction by the ground on ladder( Fr)..so, Fr=246.7 NAnd N2=900+750=1650 N Fr=u*N2u=246.7/1650=0.15Ie..the coefficient of friction between ladder and ground is 0.15